OpenStudy (anonymous):

Plot the stationary and inflection points for f(x)= x^4-3x^3+3x^2+1

OpenStudy (amorfide):

To get inflexion point you want to differentiate the function twice can you do that?

OpenStudy (anonymous):

Yeah I'm just having trouble getting the points after I differentiate he first time

OpenStudy (amorfide):

Okay so if you differentiate 4x^3-9x^2+6x=0 make it equal to 0 to get the stationary points take our the highest factor

OpenStudy (amorfide):

x(4x^2-9x+6)=0 x=0 or 4x^2-9x+6=0

OpenStudy (amorfide):

solve for x

OpenStudy (anonymous):

Okay I'll try that :)

OpenStudy (anonymous):

So the answer I get is 9 plus or minus sqrt of -15all divided by 8

OpenStudy (amorfide):

that is correct, so there is one real solution, and 2 complex solutions, ignore the complex, so it starts when x=0, y=1 that is your coordinate for minimum point which should be a stationary point.